6. Graph the function f(x)= 4(x - 2) x(x+3) labeling all intercepts and asymptotes.
Answer #1

2 4 (0-2) 2 PO given fer) - 2(x+3) x-intercept -f()=0 G=Hoy = 0 = 4 (x-2) 2 8x*x+3) - - 4 (0-2) ? 2 y- intercept: - x=0 1-20